# Improving the prediction of individual engagement in recommendations using cognitive models

### v1

@@ 3.2. Simulation Procedure
The first 25 weeks of each mothers’ trajectory were designated as training data. Following the approach of previous studies, to train the LSTM model, the training data was reconstructed into sliding windows of 7 consecutive time steps. Then an LSTM model was trained in the entire training dataset (see Figure  1b for a visual description of the differences between the IBL and LSTM training / testing procedure).
For training of the IBL models, an IBL model was trained per mother. Each IBL model traced a mother’s trajectory through each time step up to week 25. The best-fitting profile of attribute weights per mother was determined by performing a grid search for the combination that minimized the weighted sum of one-step prediction errors (loss). The grid was constructed by taking all possible combinations of parameter values in the range (0, 0) to (5, 5) in intervals of 0.5.
If an IBL model with a specific combination of attribute weights is an adequate model of an individual’s behavioral dynamics, then the model should increase its prediction accuracy as it acquires more instances and becomes more aligned in its history with that of the modeled individual. We account for this assumption in the model fitting procedure by weighting earlier prediction errors (e.g., at weeks 1 and 2) less than later ones (e.g., at weeks 24 and 25) according to the following equations:
@@
The trained models were then tasked with generating predictions for the next 14 weeks. For the one-step prediction task, the models iteratively generated predictions for the next time step (i.e., for week 26, then for week 27, then for week 28, etc.). Consistent with the fact that the sampled mothers did not receive additional interventions beyond the first 14 weeks of their tenure, we simulated these predictions in the absence of additional simulated interventions, which then allowed us to compare the models’ engagement predictions with the ground truth data.
@@
In order to simulate the effect of an intervention on a beneficiary’s engagement level when they did not receive one in reality, it was necessary to develop a counterfactual generator. We followed the approach adopted in previous studies and trained a separate LSTM model on trajectories from the entire dataset (approximately 12000 beneficiaries), which is far larger than the subset used for our policy simulations. This counterfactual generator was used only when a mother’s simulated trajectory deviated from her ground-truth trajectory (i.e., received a simulated intervention when she did not receive one in reality or not receiving a simulated intervention when she did receive one in reality).
@@ 4.3. Elucidating individual differences with IBL parameters
Beyond more accurate predictions of engagement, another advantage of the proposed IBL approach over more general TSFs (as exemplified by the LSTM approach) is the ability to identify individual differences in beneficiaries’ behavioral characteristics. Because each beneficiary is modeled by a separate IBL model, the parameter values of each model provide potential insights into how beneficiaries differ from each other in their engagement and intervention dynamics.
@@ 5. Conclusions and Future work
In programs with limited resources to provide interventions to a large number of beneficiaries, a key question is to decide which subset of beneficiaries should receive an intervention. In this paper, we develop a novel approach that uses computational cognitive models based on Instance-based Learning Theory (Gonzalez et al., 2003; Gonzalez and Dutt, 2011) to model the human dynamic transitions of being engaged or disengaged in a program. Given the non-Markovian nature of the transitions of a beneficiary between states (Danassis et al., 2023), a cognitive algorithm that reflects the effects of time, attention, and context similarity on human memory and decisions would simultaneously provide more accurate predictions of beneficiaries’ engagement dynamics and more meaningful individual differences that can also be leveraged by other prediction methods.
@@
Future work could explore personalizing interventions and intervention schedules based on these identified individual differences in behavioral characteristics. For example, intervention-sensitive beneficiaries could be allocated more frequent but shorter calls, while transition-consistent beneficiaries could be allocated fewer but longer calls to ensure that their engagement level is maximized on each call.
One limitation of this work is the lack of an unbiased counterfactual generator to simulate the effect of recommended interventions; in our work, we used a LSTM model trained on the full dataset to generate counterfactuals. It is likely that the sequence of engagement states generated by this LSTM model does not reflect the true engagement dynamics of the actual mothers, and that the generated counterfactual states would be better predicted by another LSTM model due to the similarities in model architecture. This possibly explains the apparent contradiction between the IBL models outperforming the LSTM models on prediction accuracy (which does not involve counterfactuals) but underperforming on number of engaged mothers (which requires generating counterfactuals). Future work should prioritize testing our method with other datasets, particularly with a paradigm where the effects of recommended interventions on participants’ task engagement can be experimentally measured, rather than relying on potentially biased counterfactuals.
Another limitation is that our method currently makes predictions only from the 26th week of a mother’s pregnancy onward, as the first 25 weeks are needed to train and personalize the IBL model. A possible extension of this work could involve exploring hybrid methods, such as initially relying on predictions from other trained IBL models and gradually transitioning to the personalized model.
@@ A.1. Secondary Analysis
Our experiment falls into the category of secondary analysis of the data shared by ARMMAN. This paper does not involve the deployment of the proposed algorithm or any other baselines to the service call program. As noted earlier, the experiments are secondary analysis with approval from the ARMMAN ethics board.

### v2

@@ 3.2. Simulation Procedure
We designated the first 25 weeks of each mother’s trajectory as training data. Following the approach of previous studies, we reconstructed the training data into sliding windows of 7 consecutive time steps to train the LSTM model. Then, we trained an LSTM model on the entire training dataset (see Figure  1b for a visual description of the differences between the IBL and LSTM training / testing procedure).
Each IBL model (one for each mother) traced a mother’s trajectory through each time step up to week 25. To find the best-fitting profile of attribute weights per mother, a grid search looked for the combination that minimized the weighted sum of one-step prediction errors (loss). The grid iterates all possible combinations of parameter values in the range (0, 0) to (5, 5) in intervals of 0.5.
An IBL model should improve its predictive accuracy as it accumulates more instances and becomes more “familiar” with the individual it’s modeling. So, we weight earlier prediction errors (e.g., at weeks 1 and 2) less than later ones (e.g., at weeks 24 and 25) according to the following equations:
@@
The trained models were then tasked with generating predictions for the next 14 weeks. For the one-step prediction task, the models iteratively generated predictions for the next time step (i.e., for week 26, then for week 27, then for week 28, etc.). Consistent with the fact that the sampled mothers did not receive additional interventions beyond the first 14 weeks of their tenure, we simulated these predictions without additional simulated interventions, which allowed us to compare the models’ engagement predictions with the ground truth data.
@@
To simulate the effect of an intervention on a beneficiary’s engagement level when they did not receive one in reality, it was necessary to develop a counterfactual generator. We followed the approach adopted in previous studies and trained a separate LSTM model on trajectories from the entire dataset (approximately 12000 beneficiaries), which is far larger than the subset used for our policy simulations. This counterfactual generator only operates when a mother’s simulated trajectory deviates from her ground-truth trajectory (i.e., received a simulated intervention when she did not receive one in reality or not receiving a simulated intervention when she did receive one in reality).
@@ 4.3. Elucidating individual differences with IBL parameters
Beyond more accurate predictions of engagement, another advantage of the proposed IBL approach over more general TSFs (as exemplified by the LSTM approach) is the ability to identify individual differences in beneficiaries’ behavioral characteristics. Because a separate IBL model models each beneficiary, the parameter values of each model provide potential insights into how beneficiaries differ from each other in their engagement and intervention dynamics.
@@ 5. Conclusions and Future work
In programs with limited resources to provide interventions to a large number of beneficiaries, a key question is to decide which subset of beneficiaries should receive an intervention. In this paper, we develop a novel approach that uses computational cognitive models based on Instance-based Learning Theory (Gonzalez et al., 2003; Gonzalez and Dutt, 2011) to model the human dynamic transitions of being engaged or disengaged in a program. Given the non-Markovian nature of the transitions of a beneficiary between states (Danassis et al., 2023), a cognitive algorithm that reflects the effects of time, attention, and context similarity on human memory and decisions would simultaneously provide more accurate predictions of beneficiaries’ engagement dynamics and more meaningful individual differences that other prediction methods can also leverage.
@@
Future work could explore personalizing interventions and intervention schedules based on these identified individual differences in behavioral characteristics. For example, intervention-sensitive beneficiaries could receive more frequent but shorter calls. In comparison, transition-consistent beneficiaries could receive fewer but longer calls to maximize their engagement level on each call.
One limitation of this work is the lack of an unbiased counterfactual generator to simulate the effect of recommended interventions; in our work, we used a LSTM model trained on the full dataset to generate counterfactuals. The sequence of engagement states generated by this LSTM model most likely does not reflect the true engagement dynamics of the actual mothers, and another LSTM model would better predict the generated counterfactual states due to the similarities in model architecture. This possibly explains the apparent contradiction between the IBL models outperforming the LSTM models on prediction accuracy (which does not involve counterfactuals) but underperforming on number of engaged mothers (which requires generating counterfactuals). Future work should prioritize testing our method with other datasets, particularly with a paradigm where the effects of recommended interventions on participants’ task engagement can be experimentally measured, rather than relying on potentially biased counterfactuals.
Another limitation is that our method makes predictions only from the 26th week of a mother’s pregnancy onward, as the first 25 weeks are needed to train and personalize the IBL model. A possible extension of this work could involve exploring hybrid methods, such as initially relying on predictions from other trained IBL models and gradually transitioning to the personalized model.
@@ A.1. Secondary Analysis
Our experiment falls into the category of secondary analysis of the data shared by ARMMAN. This paper does not involve deploying the proposed algorithm or any other baselines to the service call program. As noted earlier, the experiments are secondary analysis with approval from the ARMMAN ethics board.